# Data-Driven LQR with Finite-Time Experiments via Extremum-Seeking Policy Iteration

### v1

@@ V-C Proof of Theorem 2
The proof relies on the application of Theorem 1 (cf. Section II) to system (10). Then, in order to apply Theorem 1, we need to (i) choose the design parameters $c_{1},\rho>0$ bounding the initial and final values of $V$, respectively, and (ii) satisfy the conditions required by Assumptions 12, and 3. By [50, Lemma 3.8], we recall that there exists $\psi>0$ such that
@@ VII Conclusions
Then, let us consider the function $V$ (cf. (18)) and evaluate its increment $\Delta V(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}):=V((1-\gamma)\tilde{\mathrm{z}}_{\textsc{av}}+g^{\delta}(K_{\textsc{av}}),K_{\textsc{av}}-\gamma G(K_{\textsc{av}})+\gamma p^{\delta,T}(K_{\textsc{av}}))-V(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ along the trajectories of (51). By using the bounds (49)-(57) and the definition of $\tilde{V}$ (cf. (58)), we get
$\Delta V(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$
@@
$\Delta V(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ $\leq-\gamma\left(\frac{\tilde{\nu}}{2\lambda}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+\tilde{\eta}\left\|G(K_{\textsc{av}})\right\|^{2}\right)$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}-vV(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}),$ (62)
where in $(a)$ we use the definition of $V$ (cf. (18)) and $v:=\max\{\tilde{\nu},\tilde{\eta}/\mu\}$. Then, by using (62), we can further bound (61) according to
$\Delta V(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\leq-\gamma vV(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})+\tilde{V}(\bar{\gamma}_{1},\delta,\lambda).$ (63)
@@
$\Delta V(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ $\leq-\gamma\tilde{v}V(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}),$ (64)
for all $(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\in\mathbb{R}\times\mathcal{K}$ such that $V(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\geq\rho_{\textsc{av}}$. Although (64) seems to conclude the proof, we recall that it has been obtained by assuming $K_{\textsc{av}}\in\tilde{\Omega}_{c_{0}}$. In other words, since $(\tilde{\mathrm{z}}_{\textsc{av}}^{0},K_{\textsc{av}}^{0})\in\tilde{\Omega}_{c_{0}}$ by definition of $c_{0}$, to conclude the proof we only need to prove that the set ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ is forward-invariant for system (51). To this end, consider $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\tilde{\Omega}_{c_{0}}$ and, in light of the definition of $V$ (cf. (18)), we note that
$J(K_{\textsc{av}}^{{k+1}})-J(K^{\star})$ $\leq V(\tilde{\mathrm{z}}_{\textsc{av}}^{k+1},K_{\textsc{av}}^{{k+1}})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}V(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})$
@@
where in $(a)$ we use the fact that the right-hand side of (64) is non-positive for all $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\mathbb{R}\times\mathcal{K}$ such that $V(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\geq\rho_{\textsc{av}}$, while in $(b)$ we use the fact that $\lambda\geq 1$. By looking at the definition of $\tilde{\Omega}$ in (40), the inequality (65) proves the desired invariance property of ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ and the proof concludes.

### v2

@@ V-C Proof of Theorem 2
The proof relies on the application of Theorem 1 (cf. Section II) to system (10). Then, in order to apply Theorem 1, we need to (i) choose the design parameters $c_{1},\rho>0$ bounding the initial and final values of $V_{\lambda}$, respectively, and (ii) satisfy the conditions required by Assumptions 12, and 3. By [50, Lemma 3.8], we recall that there exists $\psi>0$ such that
@@ VII Conclusions
Then, let us consider the function $V_{\lambda}$ (cf. (18)) and evaluate its increment $\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}):=V_{\lambda}((1-\gamma)\tilde{\mathrm{z}}_{\textsc{av}}+g^{\delta}(K_{\textsc{av}}),K_{\textsc{av}}-\gamma G(K_{\textsc{av}})+\gamma p^{\delta,T}(K_{\textsc{av}}))-V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ along the trajectories of (51). By using the bounds (49)-(57) and the definition of $\tilde{V}$ (cf. (58)), we get
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$
@@
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ $\leq-\gamma\left(\frac{\tilde{\nu}}{2\lambda}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+\tilde{\eta}\left\|G(K_{\textsc{av}})\right\|^{2}\right)$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}-vV_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}),$ (62)
where in $(a)$ we use the definition of $V_{\lambda}$ (cf. (18)) and $v:=\max\{\tilde{\nu},\tilde{\eta}/\mu\}$. Then, by using (62), we can further bound (61) according to
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\leq-\gamma vV_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})+\tilde{V}(\bar{\gamma}_{1},\delta,\lambda).$ (63)
@@
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ $\leq-\gamma\tilde{v}V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}),$ (64)
for all $(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\in\mathbb{R}\times\mathcal{K}$ such that $V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\geq\rho_{\textsc{av}}$. Although (64) seems to conclude the proof, we recall that it has been obtained by assuming $K_{\textsc{av}}\in\tilde{\Omega}_{c_{0}}$. In other words, since $(\tilde{\mathrm{z}}_{\textsc{av}}^{0},K_{\textsc{av}}^{0})\in\tilde{\Omega}_{c_{0}}$ by definition of $c_{0}$, to conclude the proof we only need to prove that the set ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ is forward-invariant for system (51). To this end, consider $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\tilde{\Omega}_{c_{0}}$ and, in light of the definition of $V_{\lambda}$ (cf. (18)), we note that
$J(K_{\textsc{av}}^{{k+1}})-J(K^{\star})$ $\leq V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}}^{k+1},K_{\textsc{av}}^{{k+1}})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})$
@@
where in $(a)$ we use the fact that the right-hand side of (64) is non-positive for all $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\mathbb{R}\times\mathcal{K}$ such that $V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\geq\rho_{\textsc{av}}$, while in $(b)$ we use the fact that $\lambda\geq 1$. By looking at the definition of $\tilde{\Omega}$ in (40), the inequality (65) proves the desired invariance property of ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ and the proof concludes.